Find the common ratio of the infinite geometric series with the given sum and first term s=7/9 a1=7/10
Answer by rothauserc(4718)

We use the following formula for infinite geometric series sum
S = a(1 / (1-r)) where S is the sum, a is first term and r is the common ratio, therefore
7/9 = (7/10)(1 / (1-r))
7/9 = 7 / (10(1-r))
cross multiply fractions
70(1-r) = 63
1-r = 63/70 = 9/10
r = 1 - (9/10) = 1/10
the common ratio(r) is 1/10